# A Survey on Self-Supervised Learning for Non-Sequential Tabular Data

### v1

@@ 3.1 Learning from Masked Features
The objective for masking features of a sample enables the model to learn the sample context via partially known features, which also aligns with the analogous objective of downstream applications to predict the corresponding category/value of a sample from the given features. This alignment offers a trained encoder the knowledge of inferring from features of a given sample in the downstream tasks. Inspired by Masked Autoencoder (MAE) Pathak et al. (2016), utilizing random masking of the pixels and then reconstructing them to learn numerous visual concepts, TabTransformer Huang et al. (2020) and VIME Yoon et al. (2020) optimized pretext models by recovering an input sample from its corrupted or masked variants. Specifically, TabTransformer introduced random masking and random value replacement as the transformed function. VIME identified the masked features with the mask vector estimator and imputed the masked features from the correlated non-masked features with the feature vector estimator simultaneously; for example, if the value of a feature is very different from its correlated features, this feature is likely masked. The binary mask of VIME is randomly sampled from a Bernoulli distribution.
To further improve VIME by encouraging the encoder to generate more structured and representative embeddings, TabNet Arik and Pfister (2021) devised an attention mechanism to iteratively choose the features to be masked. Contrary to learnable masking, SEFS Lee et al. (2022) proposed a feature subset generator as the transformed function by enhancing the probability of masking highly correlated features. SwitchTab Wu et al. (2024) leverages the asymmetric encoder-decoder architecture on top of the self-supervise objective of VIME, and proposes a switching mechanism. Despite the progress, the percentage of masking presents an indecisive and case-by-case issue that every downstream task requires the percentage by empirical adjustments.
@@ 3.2 Perturbation in Latent Space
To learn the generalizable context from heterogeneous characteristics of tabular data, STUNT Nam et al. (2023) meta-learns self-generated tasks from unlabeled data, which is motivated by columns that may share correlations to the downstream labels (e.g., the feature of “occupation” can be used as a substituted label for “income” before the supervised learning stage). The transformed function masks some features of a table, which are then used for k-means clustering to generate pseudo-labels. On top of the meta-learning schema, STUNT is effective in few-shot tabular scenarios. LFR Sui et al. (2023) explored random projectors to learn from unlabeled data in the scenario of lacking knowledge to augment the data; nonetheless, they uncovered that it is likely to be inferior for the scenario that has sufficient information for augmentations.
@@ 3.3 Inherent in Pre-Trained Language Models
Taking another direction to solve the feature heterogeneity issue, applying language models as an encoder is able to empower transferred knowledge across different datasets by representing tabular data with semantic text. To address the significant challenge of transferring tabular data into a natural language format, various pre-trained language models (PLMs) have been incorporated with the NS-TD problems to leverage pre-trained knowledge from natural language corpora. Several works Dinh et al. (2022); Borisov et al. (2023); Zhang et al. (2023) have directly regarded numerical features as a string, which is intuitive but mitigates the effort of preprocessing (e.g., no need to transform category features into one-hot encoding). To force the language model to interpret numerical features, TP-BERTa Anonymous (2024a) proposed relative magnitude tokenization to transfer scalar to discrete tokens by the decision tree. To eliminate the feature order bias, GReaT Borisov et al. (2023) randomly permuted the order of features.
@@ 4 Contrastive Learning of SSL4NS-TD
Another common theme of the advancements is to learn robust representations via different views or corruptions of the same input, which is achieved by maximizing similarities between similar instances and pulling over instances that are dissimilar. With the success of generating views in computer vision (CV) and masking tokens in natural language processing (NLP) Jaiswal et al. (2020), contrastive learning has been attempted in tabular applications to learn effective and generic task-agnostic representations. Formally, the formula of contrastive learning can be generally defined as:
$L_{contrastive}=\phi(e(x_{i}),e(\hat{x_{i}})),$ (3)
@@
where $\phi$ is a similarity function that compares similarities between two encoded instances, and $\hat{x_{i}}$ denotes either a different instance to form a negative pair or a variant of the same instance to form a positive pair. The projection head $g$ may need to be applied after the output of $e$ if the objective requires self-supervised labels (e.g., supervised contrastive learning).
@@
SCARF Bahri et al. (2022) is an MLP-based framework with a two-stage learning strategy: InfoNCE van den Oord et al. (2018) contrastive pre-training and supervised fine-tuning. In the pre-training stage, the given input is corrupted with a random subset of its features, which are then replaced by a random view from the marginal distribution of the corresponding features. Subsequently, InfoNCE is applied as the similarity function to encourage the sample and the variant of the corresponding sample to be close, and the sample and the variants of the other samples to be far apart. In contrast to SCARF, STab Hajiramezanali et al. (2022) aims to introduce an augmentation-free self-supervised representation learning technique that does not require the need for negative pairs. STab encodes the input sample with two MLP-based encoders (one with an additional projection head), which are weight-sharing but have different stochastic regularization, which can be viewed as model-wise contrastive learning, and then compares the negative cosine distance as the similarity function. To learn contexts across tables with disparate columns that can be used for transfer learning, feature incremental learning, and zero-shot inference, TransTab Wang and Sun (2022) contextualizes the columns and cells in tables (e.g., gender is woman instead of using a categorized number) with Transformer encoders, and pre-trains on multiple tables with vertical-partition contrastive learning that designs variants based on column-wise splitting views. Anonymous (2024b) proposed a prototype-based tabular representation learning framework to learn disentangled representations around global data prototypes, which provides global prototypes to confront similar samples while preserving original distinct information with the diversifying constraint in the latent space.
@@ 5 Hybrid Learning of SSL4NS-TD
As predictive learning and contrastive learning of SSL4NS-TD have their own unique advantages and incorporate distinct self-supervision signals, an important learning strategy is to integrate both dimensions of SSL4NS-TD into a single model to provide multifaceted self-supervised tasks. Typically, models equipped with hybrid learning require multiple projection heads for different pretext tasks, which are employed in parallel to enhance self-supervision robustness. Formally, the loss of hybrid learning can be defined as:
@@ 5.1 Perturbation + Contrastive Learning
Perturbation with contrastive learning provides a natural benefit that is able to learn robust representations without specifying the explicit knowledge of tables and captures contextualized relations between rows, columns, and even cells. SubTab Ucar et al. (2021) divides tabular data into multiple subsets with potentially overlapping columns as different views for contrastive loss and distance loss, both of which enable the model to move the corresponding samples in subsets closer to each other. To perturb features unequally for feature reconstructions, SubTab adds Gaussian noise to 1) random columns, 2) a random region of neighboring columns, or 3) random features in a sample with a binomial mask. To prevent similar features from weighing too significantly in the reconstruction loss, Chen et al. (2023) integrated a regularization matrix with the reconstruction loss. With the cooperation of classification labels, they adopted different views for contrastive learning based on the labels to maximize the similarity with the same categories and leveraged semi-supervised learning to jointly pre-train the Transformer model.
@@
In addition to the achievements of employing Transformers for tabular data, researchers have started to frame NS-TD as tokens, which have been widely used in NLP and CV domains. Several variants have been proposed to capture fine-grained representations in tabular data (e.g., cells, numerical, categories) Somepalli et al. (2022); Yang et al. (2023); Zhu et al. (2023). The major advantage is that representations can be shared across various tabular datasets and can be modeled with self-supervision. SAINT Somepalli et al. (2022) described a sample with a sequence composed of the corresponding categorical or numerical features, with a special token [CLS] appended at first, similar to BERT Devlin et al. (2019). To model invariant fine-grained feature representations from other similar samples, SAINT embeds categorical as well as numerical features and encodes them with intersample attention across different rows to pre-train on a reconstruction loss and InfoNCE contrastive loss with augmentations from the embedding space.
@@
In contrast to most existing works that perform pre-training and fine-tuning per downstream dataset, another important aspect is to pre-train tabular transformers across diverse collection tables that vary in the number and types of columns. It provides the ability to serve as foundation models on a wide range of downstream tabular applications such as ChatGPT in NLP Zhou et al. (2023). XTab Zhu et al. (2023) is a general tabular transformer pre-training on the large diversity cross-tables, which is flexible to leverage existing encoder backbones (e.g, Gorishniy et al. (2021); Somepalli et al. (2022)) and existing self-supervised strategies (e.g., reconstruction loss and contrastive loss). UniTabE Yang et al. (2023) pre-trains on large-scale (13 billion examples) tabular datasets across diverse domains with a Transformer encoder-decoder architecture. The decoder takes free-form and task-specific prompts and contextualized representations from the encoder to adaptively reason on task-specific customizations. The pre-training objective of UniTabE includes multi-cell-masking to reconstruct a portion of cells of a sample and contrastive learning to treat subsets of the same sample as positive pairs and subsets of different samples as negative pairs.
@@ 5.2 Masking + Contrastive Learning
Another contribution to the realm of hybrid learning is to combine feature masking with contrastive learning since it combines the advantages of aligning the objectives between upstream and downstream data while preserving the task-agnostic learning strategy. Compared with perturbations which leave partial information for the targeted features, masking strategies completely remove the targeted features that do not contain any original information. To accommodate different features between upstream and downstream tabular data, Levin et al. (2023) introduced a pseudo-feature approach on top of existing deep tabular models for pre-training, which is able to predict missing features in upstream data but which are present in downstream data, and leverages a contrastive pre-training strategy, similar to Somepalli et al. (2022). Ye et al. (2023) pre-trained a Transformer encoder with 2k high-quality cross-table datasets with masked table modeling to learn underlying relations between features and supervised contrastive learning to cluster samples with the same label. The insights from analyses uncover that pre-training provides more transferability over tree-based baselines. Instead of feature-agnostic SSL approaches, the motivation of DoRA Du et al. (2023) is to design a pretext task based on domain knowledge in the financial domain. They introduced an intra-sample pretext task by selecting the domain-specific feature of a sample as the self-supervised label during the pre-training stage. Inter-sample contrastive learning is also adopted based on contrastive learning to separate dissimilar samples based on the domain-specific feature.
@@ 6 Tackling Application Issues of SSL4NS-TD
Benchmark Task Type #Datasets #Samples #Features Reference
MLPCBench Classification 40 400 - 400,000 4 - 2,001 Kadra et al. (2021)
DLBench Classification + Regression 11 7,000 - 1,000,000 10 - 2,000 Shwartz-Ziv and Armon (2022)
TabularBench Classification + Regression 45 3,000 - 10,000 5 - 613 Grinsztajn et al. (2022)
TabZilla Classification 36 300 - 1,000,000 7 - 4,297 McElfresh et al. (2023)
TabPretNet Unlabeled 1,000 / / Ye et al. (2023)
Classification + Regression 1,000 540 - 48,842 6 - 81 Ye et al. (2023)
Table 2: Existing NS-TD benchmarks and a pre-trained dataset. TabPretNet consists of the unlabeled set for pre-training and the labeled set for downstream evaluations. As the unlabeled set is inaccessible, we report / as not applicable for the numbers of samples and features.
The advancement of SSL4NS-TD is highly application-oriented since tabular data represent ubiquitous practical utility in diverse domains, including medicine, finance, and many other areas McElfresh et al. (2023), and research-oriented as it remains a hurdle to explore relations between rows (sample), columns (features), as well as tables (tasks). However, the main bottlenecks of existing deep learning and machine learning models for the NS-TD applications require a large number of annotated labels and suffer from the generalization from seen to new scenarios. As SSL4NS-TD manifests effective performance by learning pretext tasks from unlabeled tabular data, we explore several emerging and prevalent applications of SSL4NS-TD, showcasing their potential in the following.
@@ 6.1 Automatic Data Engineering
Although deep learning models alleviate the burden of feature engineering compared with machine learning models, the stable performance in various tasks remains a challenge due to imbalanced, missing, and noisy data Grinsztajn et al. (2022). Huang et al. (2020) demonstrated that SSL4NS-TD has the potential to maintain robust performance in terms of these scenarios across different datasets, resulting in a reduction in manual costs with minimal engineering efforts. Lee et al. (2022) effectively utilized gate vector estimation to self-supervise the selection process of correlated features. Therefore, leveraging the capabilities of SSL4NS-TD can yield significant benefits in various data engineering applications.

### v3

@@ 3.1 Learning from Masked Features
The objective for masking features of a sample enables the model to learn the sample context via partially known features, which also aligns with the analogous objective of downstream applications to predict the corresponding category/value of a sample from the given features. This alignment offers a trained encoder the knowledge of inferring from features of a given sample in the downstream tasks. Inspired by Masked Autoencoder (MAE) [32], utilizing random masking of the pixels and then reconstructing them to learn numerous visual concepts, TabTransformer [20] and VIME [50] optimized pretext models by recovering an input sample from its corrupted or masked variants, building a framework to generalize to all tabular data. Specifically, TabTransformer introduced random masking and random value replacement as the transformed function. VIME identified the masked features with the mask vector estimator and simultaneously imputed the masked features from the correlated non-masked features with the feature vector estimator; for example, if the value of a feature is very different from its correlated features, this feature is likely masked. The binary mask of VIME is randomly sampled from a Bernoulli distribution.
To further improve VIME by encouraging the encoder to generate more structured and representative embeddings, TabNet [1] devised an attention mechanism to iteratively choose the features to be masked, enabling interpretability of the deep learning model. Contrary to learnable masking, SEFS [23] proposed a feature subset generator as the transformed function by enhancing the probability of masking highly correlated features. SwitchTab [43] leverages the asymmetric encoder-decoder architecture on top of the self-supervised objective of VIME, and proposes a switching mechanism to decouple two samples, each of which consists of mutual information that is switchable between samples and salient information that is unique to each sample. Therefore, the contrastive goal is to recover the input features from both mutual and salient information. Despite the progress, the percentage of masking presents an indecisive and case-by-case issue that every downstream task requires the percentage by empirical adjustments.
@@ 3.2 Perturbation in Latent Space
To learn the generalizable context from heterogeneous characteristics of tabular data, STUNT [30] meta-learns self-generated tasks from unlabeled data, which is motivated by columns that may share correlations to the downstream labels (e.g., the feature of “occupation” can be used as a substituted label for “income” before the supervised learning stage). The transformed function masks some features of a table, which are then used for k-means clustering to generate pseudo-labels. On top of the meta-learning schema, STUNT is effective in few-shot tabular scenarios. LFR [36] explored random projectors to learn from unlabeled data in the scenario of lacking knowledge to augment the data to serve as a universal framework that can incorporate various data modalities and domains; nonetheless, they uncovered that it is likely to be inferior for the scenario that has sufficient information for augmentations.
@@ 3.3 Inherent in Pre-Trained Language Models
Taking another direction to solve the feature heterogeneity issue, applying language models as an encoder is able to empower transferred knowledge across different datasets by representing tabular data with semantic text. To address the significant challenge of transferring tabular data into a natural language format, various pre-trained language models (PLMs) have been incorporated with the NS-TD problems to leverage pre-trained knowledge from natural language corpora. Several works [12, 8, 51] have directly regarded numerical features as a string, which is intuitive but mitigates the effort of preprocessing (e.g., there is no need to transform category features into one-hot encoding). To force the language model to interpret numerical features, TP-BERTa [46] proposed relative magnitude tokenization to transfer scalar to discrete tokens by binning numerical values with decision trees. Subsequently, the embeddings of binned tokens are then multiplied by the original values to avoid a large number of values within a single bin. To eliminate the feature order bias, GReaT [8] connects tabular and text modalities with a textual encoding schema and randomly permutes the order of features. For instance, the schema converts the input tabular features “age = 26; income = 70k” as “age is 26, income is 70k” to be consumed by LLMs. The order perturbation is then perturbed as “income is 70k, age is 26” as synthetic data generation results.
@@ 4 Contrastive Learning of SSL4NS-TD
Another common theme of the advancements is to learn robust representations via different views or corruptions of the same input, which is achieved by maximizing similarities between similar instances and pulling over instances that are dissimilar. With the success of generating views in computer vision (CV) and masking tokens in natural language processing (NLP) [21], contrastive learning has been attempted in tabular applications to learn effective and generic task-agnostic representations. Formally, the formula of contrastive learning can be generally defined as:
$L_{contrastive}=\phi(h_{i},h_{i}^{+},h_{j}^{-}),$ (3)
$h_{i}=e(x_{i});h_{i}^{+}=e(x_{i}^{+});h_{j}^{-}=e(x_{j}^{-}),$ (4)
@@
where $\phi$ is a similarity function that compares similarities between the anchor ($h_{i}$), positive ($h_{i}^{+}$), and negative ($h_{i}^{-}$) instances. A positive instance is a variant of the anchor by augmentation methods (e.g., partial features of the anchor), and a negative instance is often sampled from the other instances. Positive pairs represent two similar instances, while negative pairs represent two dissimilar instances. Generally, the similarity function can be opted by cosine similarity, Euclidean distance, or dot product to compute pairwise similarity. Note that positive and negative pairs may be optional depending on the contrastive loss as depicted next. The projection head $g$ may need to be applied after the output of $e$ if the objective requires self-supervised labels (e.g., supervised contrastive learning).
SCARF [4] is an MLP-based framework with a two-stage learning strategy: InfoNCE [31] contrastive pre-training and supervised fine-tuning, which reinforces the generalizability for various tabular domains. In the pre-training stage, the given input is corrupted with a random subset of its features, which are then replaced by a random view from the marginal distribution of the corresponding features to represent positive and negative instances. Subsequently, InfoNCE is applied as the similarity function to encourage the sample and the variant of the corresponding sample to be close, and the sample and the variants of the other samples to be far apart:
$\phi_{InfoNCE}(h_{i},h_{i}^{+},h_{j}^{-})=log(\frac{exp(\frac{h_{i}\cdot h_{i}^{+}}{\tau})}{\sum_{j=1}^{N}exp(\frac{h_{i}\cdot h_{j}^{-}}{\tau})}),$ (5)
where $\tau$ is the temperature parameter.
@@
In contrast to SCARF, STab [16] aims to introduce an augmentation-free self-supervised representation learning technique that does not require the need for negative pairs. Specifically, STab encodes the input sample with two MLP-based encoders (one with an additional projection head), which are weight-sharing but have different stochastic regularization, which can be viewed as model-wise contrastive learning, and then compares the negative cosine distance as the similarity function. The two representations are therefore considered as a corrupt version of the other one instead of selecting from the subset of its features. To learn contexts across tables with disparate columns that can be used for transfer learning, feature incremental learning, and zero-shot inference, TransTab [40] contextualizes the columns and cells in tables (e.g., gender is woman instead of using a categorized number) with Transformer encoders, and pre-trains on multiple tables with vertical-partition contrastive learning that designs variants based on column-wise splitting views. [49] proposed a prototype-based tabular representation learning framework to learn disentangled representations around global data prototypes, which provides global prototypes to confront similar samples while preserving original distinct information with the diversifying constraint in the latent space.
@@ 5 Hybrid Learning of SSL4NS-TD
As predictive learning and contrastive learning of SSL4NS-TD have their own unique advantages and incorporate distinct self-supervision signals, an important learning strategy is to integrate both dimensions of SSL4NS-TD into a single model to provide multifaceted self-supervised tasks. Typically, models equipped with hybrid learning require multiple projection heads for different pretext tasks (i.e., multi-tasking), which are employed in parallel to enhance self-supervision robustness. Formally, the loss of hybrid learning can be defined as:
@@ 5.1 Perturbation + Contrastive Learning
Perturbation with contrastive learning provides a natural benefit that is able to learn robust representations without specifying the explicit knowledge of tables, and captures contextualized relations between rows, columns, and even cells. In addition to the reconstruction loss, SubTab [38] divides tabular data into multiple subsets with potentially overlapping columns as different views (similar to the idea of cropping images in CV) for contrastive loss and distance loss, both of which enable the model to move the corresponding samples in subsets closer to each other. To perturb features unequally for feature reconstructions, SubTab adds Gaussian noise to 1) random columns, 2) a random region of neighboring columns, or 3) random features in a sample with a binomial mask. To prevent similar features from weighing too significantly in the reconstruction loss, [9] integrated a regularization matrix with the reconstruction loss. With the cooperation of classification labels, they adopted different views for supervised contrastive learning based on the labels to maximize the similarity with the same categories, and leveraged semi-supervised learning to jointly pre-train the Transformer model.
@@
In addition to the achievements of employing Transformers for tabular data, researchers have started to frame NS-TD as tokens, which have been widely used in NLP and CV domains. Several variants have been proposed to capture fine-grained representations in tabular data (e.g., cells, numerical, categories) [35, 53, 47]. The major advantage is that representations can be shared across various tabular datasets and can be modeled with self-supervision. SAINT [35] described a sample with a sequence composed of the corresponding categorical or numerical features, with a special token [CLS] appended at first, similar to BERT [11]. To model invariant fine-grained feature representations from other similar samples, SAINT embeds categorical as well as numerical features and encodes them with intersample attention across different rows to pre-train on a reconstruction loss and InfoNCE contrastive loss with augmentations from the embedding space.
@@
In contrast to most existing works that perform pre-training and fine-tuning per downstream dataset, another important aspect is to pre-train tabular transformers across diverse collection tables that vary in the number and types of columns. It provides the ability to serve as foundation models on a wide range of downstream tabular applications such as ChatGPT in NLP [52]. XTab [53] is a general tabular transformer pre-training on the large diversity cross-tables, which is flexible to leverage existing encoder backbones (e.g., [14, 35]) and existing self-supervised strategies (e.g., reconstruction loss and contrastive loss). UniTabE [47] pre-trains on large-scale (13 billion examples) tabular datasets across diverse domains with a Transformer encoder-decoder architecture. The decoder takes free-form and task-specific prompts as well as contextualized representations from the encoder to adaptively reason on task-specific customizations. That is, the prompt can be modified to adapt to specific downstream applications. The pre-training objective of UniTabE includes multi-cell-masking to reconstruct a portion of cells of a sample and contrastive learning to treat subsets of the same sample as positive pairs and subsets of different samples as negative pairs.
@@ 5.2 Masking + Contrastive Learning
Another contribution to the realm of hybrid learning is to combine feature masking with contrastive learning since it combines the advantages of aligning the objectives between upstream and downstream data while preserving the task-agnostic learning strategy. Compared with perturbations which leave partial information for the targeted features, masking strategies completely remove the targeted features that do not contain any original information. To accommodate different features between upstream and downstream tabular data, [25] introduced a pseudo-feature approach on top of existing deep tabular models for pre-training, which is able to predict missing features in upstream data but which are present in downstream data, and leverages a contrastive pre-training strategy, similar to [35]. [48] pre-trained a Transformer encoder with 2k high-quality cross-table datasets with masked table modeling to learn underlying relations between features and supervised contrastive learning to cluster samples with the same label. The insights from analyses uncover that pre-training provides more transferability over tree-based baselines. Instead of feature-agnostic SSL approaches, the motivation of DoRA [13] focuses on designing a pretext task based on domain knowledge in the financial domain for real estate appraisal. They introduced an intra-sample pretext task by selecting the domain-specific feature of a sample as the self-supervised label during the pre-training stage (i.e., predict the located town of the given real estate). Inter-sample contrastive learning is also adopted based on contrastive learning to separate dissimilar samples based on the domain-specific feature (i.e., two real estates located in the same town are closer).
@@ 6 Tackling Application Issues of SSL4NS-TD
Benchmark Task Type #Datasets #Samples #Features
OpenML-CC18[1] C 72 500 - 92,000 5 - 3,073
DLBench[2] C + R 11 7,000 - 1,000,000 10 - 2,000
TabularBench[3] C + R 45 3,000 - 10,000 5 - 613
TabZilla[4] C 36 300 - 1,000,000 7 - 4,297
TP-BERTa[5] U 202 10,000 - 100,000 1 - 31
C + R 145 10 - 9,800 3 - 32
OpenTabs[6] U 2,000 23,000* 24*
UniTabE[7] U 283,000 46,000* 31*
[1][7], [2][34], [3][15], [4][29], [5][46], [6][48], [7][47]
Table 2: Existing NS-TD benchmarks and a pre-trained dataset. TabPretNet consists of the unlabeled set for pre-training and the labeled set for downstream evaluations. “C” denotes classification, “R” denotes regression, and “U” denotes unlabeled. “*” indicates the numbers are only provided the average numbers in their original papers.
The advancement of SSL4NS-TD is highly application-oriented since tabular data represent ubiquitous practical utility in diverse domains, including medicine, finance, and many other areas [29], and research-oriented as it remains a hurdle to explore relations between rows (sample), columns (features), as well as tables (tasks). However, the main bottlenecks of existing deep learning and machine learning models for the NS-TD applications demand human engineering efforts, require a large number of annotated labels, and struggle with generalization from known to new scenarios. As SSL4NS-TD manifests effective performance by learning pretext tasks from unlabeled tabular data, we explore several emerging and prevalent applications of SSL4NS-TD, showcasing their potential in the following.
@@ 6.1 Automatic Data Engineering
Although deep learning models alleviate the burden of feature engineering compared with machine learning models, the stable performance in various tasks remains a challenge due to imbalanced, missing, and noisy data [15]. [20] demonstrated that SSL4NS-TD has the potential to maintain robust performance in terms of these scenarios across different datasets, resulting in a reduction in manual costs with minimal engineering efforts. [23] effectively utilized gate vector estimation to self-supervise the selection process of correlated features that explicitly avoid selecting redundant features and allow informative features to be learned. Therefore, leveraging the capabilities of SSL4NS-TD can yield significant benefits in various data engineering applications.